OpenStudy (roseflight):

Having some trouble with exponential problems... Ok, so say you've got e ^(2x)=e ^(x+1). How would you go about finding a value for x? Alternatively, if you had 4^(-x)=2^(x+1) how would you find THAT x? Any help with either or both of those problems would be extremely useful :)

OpenStudy (anonymous):

On taking log on the both sides of the Eqn \[e ^{2x}=e ^{x+1} \] become 2x=x+1 Hence x=1 is the answer

OpenStudy (anonymous):

Apply same method for the other question